OpenStudy (iwanttogotostanford):

HELP PLEASE: Verify the Identity cot[x-pi/2] = -tanx

OpenStudy (evoker):

Cot(A-B) = (CotACotB+1)/(CotB-CotA), you will probably need that identity

OpenStudy (bobo-i-bo):

Do you know the definition of \(\cot(\theta)\)?

OpenStudy (phi):

you could change everything to sin and cos and then use sin(A-B)= sin A cos B - cos A cos B and cos(A-B) = cos A cos B + sin A sin B and also remember sin pi/2 = 1 cos pi/2 = 0